OpenStudy (anonymous):

Please helppp!! Given the equation Square root of 2x plus 1 = 3, solve for x and identify if it is an extraneous solution. x = 4, solution is extraneous x = 4, solution is not extraneous x = 5, solution is extraneous x = 5, solution is not extraneous

OpenStudy (camerondoherty):

\[\sqrt{2x}+1=3\]

OpenStudy (abmon98):

\[\sqrt{2x+1}=3\] \[3^2=2x+1\] \[2x=8\] \[x=4\]

OpenStudy (camerondoherty):

so its B because its not extraneous

OpenStudy (camerondoherty):

but give medal to @Abmon98

OpenStudy (camerondoherty):

I take no credit

OpenStudy (abmon98):

@camerondoherty To be honest, I dont know what extraneous means.

OpenStudy (camerondoherty):

extraneous means basically its FALSE and non extraneous means its TRUE with the solution i guess if you narrow it down

OpenStudy (camerondoherty):

thats what my teacher told me and how i learned it

OpenStudy (anonymous):

Thanks ! ♥

OpenStudy (anonymous):

can any of you help me with one more question?

OpenStudy (camerondoherty):

ask a new question and tag me in it like this @camerondoherty

OpenStudy (anonymous):

Eloise started to solve a radical equation in this way: Square root of negative 2x plus 1 − 3 = x Square root of negative 2x plus 1 − 3 + 3 = x + 3 Square root of negative 2x plus 1 = x + 3 Square root of negative 2x plus 1 − 1 = x + 3 − 1 Square root of negative 2 x = x + 2 (Square root of negative 2 x)2 = (x − 4)2 −2x = x2 − 8x + 16 −2x + 2x = x2 + 8x + 16 + 2x 0 = x2 + 10x + 16 0 = (x + 2)(x + 8) x + 2 = 0 x + 8 = 0 x + 2 − 2 = 0 − 2 x + 8 − 8 = 0 − 8 x = −2 x = −8 Both solutions are extraneous because they don't satisfy the original equation. What error did Eloise make? She added 2x after squaring both sides. She subtracted 1 before squaring both sides. She factored x2 + 10x + 16 incorrectly. She did not check for extraneous solutions.

OpenStudy (anonymous):

@camerondoherty idk how to take a picture of it so i had to put it here :/

OpenStudy (camerondoherty):

can you link me the page ill take a puctiure

OpenStudy (anonymous):

its on FLVS..

OpenStudy (camerondoherty):

because i cant really understand that xD

OpenStudy (camerondoherty):

just try it

OpenStudy (anonymous):

lol okay let me see if i can hold on

OpenStudy (anonymous):

@camerondoherty

OpenStudy (camerondoherty):

oh gawsh idk sorry...

OpenStudy (anonymous):

its okay thanks anyways !! :D